OpenStudy (anonymous):

help: What is the general form of the equation for the given circle? x2 + y2 − 8x − 8y + 23 = 0 x2 + y2 − 8x − 8y + 32 = 0 x2 + y2 − 4x − 4y + 23 = 0 x2 + y2 + 4x + 4y + 9 = 0

OpenStudy (anonymous):

@undeadknight26 @Deadly_Roses @GabeBae

OpenStudy (gabebae):

What is this?

OpenStudy (anonymous):

hmmm. is this even math?

OpenStudy (anonymous):

@GabeBae

OpenStudy (anonymous):

yes its a Math question that I don't get at all

OpenStudy (anonymous):

If this has something to do with circles. What grade are you?

OpenStudy (anonymous):

if you could answer that?

OpenStudy (zpupster):

ok do you know how to complete a square?

OpenStudy (zpupster):

(x^2 -8x _) + y^2- 8y _) +23=0

OpenStudy (zpupster):

take the first equation in your answers and separate the terms like this (x^2 -8x _) + y^2- 8y _) +23=0 now half the middle term and sqaure it like this 1/2)(-8) = -4 and (-4)^2 = 16, x and y are the same, remember we are adding 16 onthe left twice--so we must add it on the right twice,so (x^2 -8x+16) + (y^2- 8y+16) +23=16+16 now factor the x nad y trinomial into squares what do you get??

OpenStudy (zpupster):

can you factor them??